Status Report

NASA ISS On-Orbit Status 14 March 2009

By SpaceRef Editor
March 14, 2009
Filed under , , ,
NASA ISS On-Orbit Status 14 March 2009
http://images.spaceref.com/news/iss.66.jpg

All ISS systems continue to function nominally, except those noted previously or below. Saturday – half-day off for CDR Fincke, FE-1 Lonchakov & FE-2 Magnus. >>>Today is Commander Michael Fincke’s 42nd birthday. Happy Birthday, Mike!<<< For math freaks worldwide, 3/14 is also “Happy Pi Day” (3.1415926…).

With STS-119/15A launch re-scheduled for not earlier than tomorrow evening, the crew’s wake period is again shifting back from yesterday, to 7:30am-9:00pm EDT, to adjust to a Discovery arrival on 3/17 and the subsequent docked period.

In the DC1 (Docking Compartment), FE-1 Lonchakov terminated the discharge process on the second 825M3 Orlan battery pack in the ZU-S recharge unit.

Afterwards, the FE-1 also spent another two hours on post-EVA clean-up operations, stowing tools while keeping track of the moves in the IMS (Inventory Management System). Ground support was available for tagup as required.

At ~9:20am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

CDR Fincke performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The station residents completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1/2.5h) and ARED advanced resistive exerciser (CDR, FE-2).

Later today, FE-2 Magnus is scheduled to terminate the recharge of REBA (Rechargeable EVA Battery) #1008 in the US Airlock.

Sandra will also perform the regular bi-monthly reboots of the OCA Router and File Server SSC (Station Support Computer) laptops.

The crew is timelined for the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

In the JAXA “Kibo” JPM (JEM Pressurized Module), Sandy will be assisting SSIPC/Tsukuba (Space Station Integration & Promotion Center) Flight Controllers in more troubleshooting the failed CBEF (Cell Biology Experiment Facility) Micro-G Incubator temperature controller fan. [On 2/28, the fan was found to be not working, depriving the system of essential temperature control (neither too cold nor too hot). The fan is a three-phase motor which mechanically stops at 12 different positions by magnet force of the fan motor, and the potential cause of the anomaly is breaking of one of the magnetic stoppage coils. Sandy’s support consists in rotating the fan to its non-startable position and observing/reporting whether it starts rotating this time.]

At ~2:20pm, the CDR will have his weekly PFC (Private Family Conference), via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), today clearly to receive birthday congrats from his family.

Weekly Science Update (Expedition Eighteen — Week 20)

3-D SPACE: Planned: 15A/FE-2.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): Planned.

BIO-4: Complete.

BIOLAB (BLB): Underway.

BIOPHOSPHONATES: Planned (15A FE-2)

CARDIOCOG-2: Complete.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Sandy, thank you for completing your second session of CCISS. We appreciate your patience in working the procedures piece by piece.”

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Ongoing.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): Ongoing.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

DomeGene (JAXA): “Regarding the troubleshooting on the failure of the temperature controller fan in the CBEF Micro-G Incubator, thank you for your activity on 3/8 to rotate the fan with a cotton swab. It worked, and we now know the fan starting position. We are now planning another data acquisition using ground commanding for the risk mitigation as preparation for Dome Gene experiment.”

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

EDR (European Drawer Rack): Parameter update in preparation for EDR performed on 3/11.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): ENose is operating normally.

EPM (European Physiology Module): Underway.

EPO (Educational Payload Operations): EPO-CEO-Demo was completed during Voluntary Science Time. The crew created a 42 minute video to provide students in grades 5-12 and educators with a stronger understanding of the CEO (Crew Earth Observation) photos being taken on-board the ISS. The video is also intended to introduce the Expedition Earth and Beyond Education Program. Also, EPO-Renovation-Demo was completed during Voluntary Science Time. The crew created a 53 minute video intended for students in grades K-12 and educators with a stronger understanding of how astronauts on-board the ISS are constructing and renovating the ISS to prepare for larger crews and more research in the future.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): The platform continues to operate nominally. DEBIE-2, DOSTEL, MEDET, FIPEX and EXPOSE are acquiring science. DEBIE-2 troubleshooting ongoing on the ground model to try and understand and solve the problem of the empty science packets. On FIPEX a link error occurred while running a science script on 3/8 and FIPEX was rebooted. A new science acquisition script started on 3/11. MEDET has started a campaign with a modified parameter for the detection rate of the spectrometer. Ground commanding took place on 3/10 for this 1-week measurement campaign. TRIBOLAB was commanded into Thermal Stabilization Mode on 3/12 to prepare for further attempts to resume Ball Bearing run#4.

FSL (Fluid Science Laboratory): Ongoing.

GEOFLOW: Aborted. GEOFLOW EC (Experiment Container) will be returned on 15A for investigation of its anomaly.”

HDTV System Test DL (JAXA): Complete.

ICE CRYSTAL (JAXA): Complete.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Underway.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

Micro-G Clay (JAXA EPO): Complete.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): Complete.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY: “Sandy, thank you for completing your FD-120 session. We appreciate the information on consumable numbers. This will help us in planning for Nicole’s sessions.”

PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PCDF-PU (Protein Crystallization Diagnostic Facility – Process Unit): Parameter update in preparation of PCDF performed successfully on 3/11. PCDF reactors awaiting launch.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

RadGene & LOH (JAXA): “Thank you Sandy for completing this experiment. Because of the launch delay, and also for the CO2 gas problem, we had to re-schedule so many times! But, finally we can complete the experiment. PI team is very pleased with the completion. We are waiting for the 15A launch.”

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Thank you, Mike, for completing your week of Sleep logs. This closes out your required logging.”

SOLAR (Solar Monitoring Observatory): The platform has been in Idle Mode during the last reporting period as it is out of the Sun visibility window. The next window is predicted to start on 3/22.

SOLO (Sodium Loading in Microgravity): Experiment completed by Mike Fincke. Blood and urine samples are now stowed in MELFI for return on 2J/A.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

SPICE (Smoke Point In Co-flow Experiment): “All 108 nominal test points have been completed along with an almost equal number of bonus runs. Thanks to Mike and Sandy’s efficient work. Some of the fuels still have significant amounts of gas left which will be translated into additional test points after the ongoing analysis of the completed test points is completed. The double bonus test matrix is expected to be completed in two weeks.”

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): Through 3/10, the ground has received a total of 23,842 of ISS CEO imagery for review and cataloguing. “Photos with times corresponding to our CEO target times are reviewed first and this week include: Heard Island – 2 frames – target not acquired; Patagonian Glaciers – 12 frames – under review; Ulawun Volcano – 6 frames – target not acquired; Villarrica Volcano, Chile – 39 frames – target acquired – excellent imagery – another target on our list bites the dust; South Georgia/ South Sandwich Islands – 24 frames – target acquired – but more clouds than we had hoped; and Coral Reefs – American Samoa – target acquired – under review. In addition: On your own initiative and time you completed yet another CEO target requirement – Heard Island, South Indian Ocean – we’ve never seen better views even from satellites and we’re ecstatic. In fact, one of your dramatic, low sun-lit views from this session of Mawson Peak, a volcanic cone in breached caldera of larger Big Ben volcano, will be published on NASA/GSFC’s Earth Observatory website this weekend. This target has been attempted dozens of times by ISS crews without success until now. Kudos and thanks for your diligence!”

CEO (Crew Earth Observations) photo targets uplinked for today were Patagonian Glaciers (ISS orbit passed over the Northern Patagonian Ice Field located in Chile. The San Rafael Glacier is one of the major outlet glaciers of the Northern Patagonian Ice Field in southern Chile and it calves into the Laguna San Rafael. The crew was successful in capturing the terminus of this glacier earlier in their increment. CEO researchers requested mapping of this glacier near nadir, using the 800 mm lens to follow the glacier from its end [terminus] back to its source [ice field]. The was in mid-morning with partly cloudy views expected), Patagonian Glaciers (on this second daylight-awake pass over Patagonia in late morning, partly cloudy views were expected for the extreme southern portion of this target area. CEO interest here is Tyndall Glacier. The southernmost of several long valley glaciers on the southeastern flank of the Southern Patagonian Ice Field, Tyndall calves into Geikie Lake and retreated significantly in recent years. Using the long lens settings to update CEO imagery with a mapping of the length of the glacier), Patagonian Glaciers (ISS had a third midday, near-nadir pass with at least a partly cloudy views over the southern portion of the Southern Patagonian Ice Field. Specifically CEO researchers are seeking near-nadir views of the Perito Moreno Glacier, a large [250 square kilometer in area with a length of 30km] feature on the southeastern flank of the ice field. It is one of three Patagonian glaciers that are not retreating. “Please update our imagery of this magnificent glacier”), Chaiten Volcano (ISS had a near nadir, early afternoon pass with partly cloudy weather over this recently reactivated volcano in southern Chile. Prior to its eruption in May 2008, the volcano had been quiet for more than 9,000 years; it has caused significant damage to the town of Chaiten located to the SW. Looking just left of track for the volcano; photography of the summit lava domes is of particular interest. Steam and ash plumes may also have been visible. Last month on 2/24 while Mike & Sandy successfully captured amazing imagery of Villarrica volcano they were also able to capture Chaiten with the 180 mm lens. This time the ground was asking for a tighter view with the 800 mm lens), and Lake Poopo, Bolivia (a persistent pattern of clouds and weather over the Altiplano and the Lake Poopo target area has final broken up. The crew’s last view of Lake Poopo showed it to be all but completely dry. This pass was in late afternoon light and over the extreme southeastern corner of the target area so the ground recommended short lens views of Poopo, the Coipasa and Uyuni salars, and obliques north-northwestward toward Lake Titicaca. “Let’s see if there was any discernible precipitation-increase in water levels from recent heavy cloud cover.”)

CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 8:08am EDT [= epoch])
Mean altitude — 354.7 km
Apogee height — 361.1 km
Perigee height — 348.3 km
Period — 91.63 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0009551
Solar Beta Angle — 48.1 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 64 m
Revolutions since FGB/Zarya launch (Nov. 98) — 59105

Significant Events Ahead (all dates Eastern Time, some changes possible!):
03/15/09 — STS-119/Discovery/15A launch – S6 truss segment — 7:43:40pm EDT (NET)
03/17/09 — STS-119/Discovery/15A docking — 5:13pm EDT
03/23/09 — STS-119/Discovery/15A undocking
03/25/09 — STS-119/Discovery/15A deorbit & landing
03/26/09 — Soyuz TMA-14/18S launch (7:49am EDT)
03/28/09 — Soyuz TMA-14/18S docking (SM aft port; 9:14am EDT)
04/07/09 — Soyuz TMA-13/17S undocking (1:02am) & landing (4:20am EDT)
05/06/09 — Progress 32P undocking & deorbit
05/07/09 — Progress 33P launch
05/12/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/12/09 — Progress 33P docking
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/27/09 — Soyuz TMA-15/19S launch
05/29/09 — Soyuz TMA-15/19S docking (FGB nadir)
Six-person crew on ISS
07/17/09 — Progress 33P undock & deorbit
07/20/09 — Soyuz TMA-14/18S relocation (to DC1)
07/24/09 — Progress 34P launch
07/26/09 — Progress 34P docking (SM aft)
08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC
09/01/09 — H-IIB (JAXA HTV-1) — tentative
11/10/09 — Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz — tentative
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola — tentative
02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC — tentative
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1 — tentative
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 — tentative
12/XX/11 — Proton 3R/MLM w/ERA.

SpaceRef staff editor.